# Optimistic Query Routing in Clustering-based Approximate Maximum Inner Product Search

### v3

@@ Abstract.
Clustering-based nearest neighbor search is an effective method in which points are partitioned into geometric shards to form an index, with only a few shards searched during query processing to find a set of top-$k$ vectors. Even though the search efficacy is heavily influenced by the algorithm that identifies the shards to probe, it has received little attention in the literature. This work bridges that gap by studying routing in clustering-based maximum inner product search. We unpack existing routers and notice the surprising contribution of optimism. We then take a page from the sequential decision making literature and formalize that insight following the principle of “optimism in the face of uncertainty.” In particular, we present a framework that incorporates the moments of the distribution of inner products within each shard to estimate the maximum inner product. We then present an instance of our algorithm that uses only the first two moments to reach the same accuracy as state-of-the-art routers such as Scann by probing up to $50\%$ fewer points on benchmark datasets. Our algorithm is also space-efficient: we design a sketch of the second moment whose size is independent of the number of points and requires $\mathcal{O}(1)$ vectors per shard.
maximum inner product search, clustering-based nearest neighbor search, query routing, sketching, optimism principle
ccs: Information systems Retrieval models and ranking
1. Introduction
A fundamental operation in modern information retrieval and database systems is what is known as Nearest Neighbor search or top-$k$ vector retrieval (Bruch, 2024). It is defined as follows: Given a collection $\mathcal{X}$ of $m$ data points in $\operatorname*{\mathbb{R}}^{d}$, we wish to find the $k$ closest points to a query $q\in\operatorname*{\mathbb{R}}^{d}$, where closeness is determined by some notion of vector similarity or distance. In this work, we focus exclusively on inner product as a measure of similarity, leading to the following formal definition known as Maximum Inner Product Search (MIPS):
(1) $\mathcal{S}=\operatorname*{arg\,max\quad}_{u\in\mathcal{X}}^{(k)}\langle q,u\rangle.$
@@ 1. Introduction
Unsurprisingly, it is often too difficult to solve this problem exactly within a reasonable time budget, especially as $m$ or $d$ increases. As such, the problem is often relaxed to its approximate variant, where we tolerate error in the retrieved set to allow faster query processing. This approximate version of the problem is aptly named Approximate Nearest Neighbor (ANN) search, whose effectiveness is characterized by the fraction of true nearest neighbors recalled in the retrieved set: $\lvert\tilde{\mathcal{S}}\cap\mathcal{S}\rvert/k$, where $\tilde{\mathcal{S}}$ is the set returned by an ANN algorithm. We call this metric recall, or equivalently, accuracy.
1.1. Clustering-based ANN search
@@
ANN algorithms come in various flavors, from trees (Bentley, 1975; Dasgupta and Sinha, 2015), LSH (Indyk and Motwani, 1998), to graphs (Malkov and Yashunin, 2020; Jayaram Subramanya et al., 2019). Refer to (Bruch, 2024) for a thorough review of this subject. The method relevant to this work is the clustering-based approach, also known as Inverted File (IVF) (Jégou et al., 2011), which has proven effective in practice (Auvolat et al., 2015; Babenko and Lempitsky, 2012; Chierichetti et al., 2007; Bruch et al., 2024; Douze et al., 2024). In this paradigm, data points are partitioned into $C$ shards using a clustering function $\zeta:\operatorname*{\mathbb{R}}^{d}\rightarrow[C]$ on $\mathcal{X}$. A typical choice for $\zeta$ is the KMeans algorithm with $C=\mathcal{O}(\sqrt{m})$. This forms the index data structure.
Accompanying the index is a routing function $\tau:\mathbb{R}^{d}\rightarrow[C]^{\ell}$. It takes a query $q$ and returns $\ell$ shards that are more likely to contain its nearest neighbors. A common router is as follows:
(2) $\tau(q)=\operatorname*{arg\,max\quad}^{(\ell)}_{i\in[C]}\langle q,\mu_{i}\rangle,$
where $\mu_{i}$ is the mean of the $i$-th shard.
@@
Processing a query $q$ involves two independent subroutines. The first stage, which we call “routing,” obtains a list of $\ell$ shards using $\tau(q)$, and the subsequent step, “scoring,” computes inner product between the query and the points in the union of the selected shards. While a great deal of research has focused on scoring step (Jégou et al., 2011; Ge et al., 2014; Wu et al., 2017; Andre et al., 2021; Kalantidis and Avrithis, 2014; Johnson et al., 2021; Norouzi and Fleet, 2013), routing has received relatively little attention. This work turns squarely to the first step: Routing queries to shards.
@@
Before we proceed further, we wish to emphasize the following point: In a clustering-based ANN system, we can select a routing protocol to choose shards, independently of the scoring algorithm that evaluates the chosen shards. For example, we may route queries to shards (i.e., decide which shards to probe) using Equation (2) or using our novel router to be introduced later. This choice has no algorithmic bearing on the scoring phase whatsoever. In other words, the scoring stage can perform a linear scan of the chosen shards, or it may compute inner products using Product Quantization (PQ) (Jégou et al., 2011), or it may further index and search each shard using a graph ANN algorithm (Jayaram Subramanya et al., 2019; Malkov and Yashunin, 2020). All these choices are independent of the type of router utilized in the earlier stage. This is indeed one of the advantages of a clustering-based ANN search system.
That independence simplifies the evaluation of different routing protocols: By independence and the additivity of latency, any efficiency savings achieved by a router translates directly into efficiency savings, of the same magnitude, for the end-to-end retrieval system (i.e., routing followed by scoring). For instance, if two routers lead to the same end-to-end recall, but one does so at a substantially lower latency, its latency savings will be reflected in its end-to-end query latency. As a result, we can fix the scoring stage across our experiments (to an exact linear scan, e.g.), and examine routers in isolation to compare their accuracy, latency, and other characteristics.
1.2. The importance of routing
@@
The historical focus on the scoring phase makes sense. Even though the routing step narrows down the search space, selected shards may nonetheless contain a large number of points. It is thus imperative that the scoring stage be efficient and effective.
@@
We argue that the oft-overlooked routing step is important in its own right. The first and obvious reason is that, the more accurately [^1] We quantify the accuracy of a router as the ANN recall in a setup where the scoring stage performs a linear scan over the identified shards. As such, the only source of ANN error is the routing procedure, rather than inner product computation. a router chooses a subset of shards, the fewer data points the scoring stage must examine. For example, if shards are balanced in size, access to an oracle router (i.e., one that identifies the shard with the true nearest neighbor) means that the scoring stage must only compute inner products for $m/C$ points to find the top-$1$ point.
@@
The second reason pertains to scale. As collections grow in size and dimensionality, it is often infeasible to keep the entire index in memory, in spite of compression techniques such as PQ. Much of the index must therefore rest on secondary storage—in particular, cheap but high-latency storage such as disk or blob storage—and accessed only when necessary. That line of reasoning has led to the emergence of disk-based graph indexes (Jayaram Subramanya et al., 2019; Singh et al., 2021; Jaiswal et al., 2022) and the like.
@@
Translating the same rationale to the clustering-based paradigm implies that shards rest outside of the main memory, and that when a router identifies a subset of shards, the scoring stage must fetch those shards from storage for further processing. This paradigm has gained traction in real-world vector database systems. [^2] See, for example, https://turbopuffer.com/blog/turbopuffer and https://www.pinecone.io/blog/serverless-architecture/#Pinecone-serverless-architecture. This is the context in which we present our research: Storage-backed, clustering-based ANN search systems.
@@
In this framework, a more accurate router lowers the volume of data that must be transferred between storage and memory. That has several consequences including a lower utilization of the storage bandwidth per query, as well as a lower consumption of in-memory cache space per query, thereby improving retrieval throughput.
1.3. An illustrative example
Interestingly, depending on operational factors such as query load, storage I/O bandwidth, data transfer rate, and memory utilization, it would be acceptable for the routing stage to be more computationally expensive as long as it identifies shards more accurately. Let us consider an illustrative example to support this statement and, in fact, motivate our research.
Consider Natural Questions (Kwiatkowski et al., 2019) embedded with Ada-002333https://openai.com/index/new-and-improved-embedding-model/ denoted by Nq-Ada2. It contains $2.7$ million $1{,}536$-dimension points. When clustered with KMeans, each shard contains about $1{,}650$ points, and is $1$MB in size, if compressed with $4$-bit PQ codebooks—a common configuration.
Next, consider a commercial machine: AWS c5.xlarge (with $4$ vCPUs and $8$GB of memory). Using $4$ threads, transferring $4$MB of data from blob storage (hosted on Amazon’s own S3) to main memory has an average latency of $45$ milliseconds (ms). [^4] See https://github.com/dvassallo/s3-benchmark for an independent and comprehensive benchmark.
Given this setup, fetching $16$ Nq-Ada2 shards from blob storage and downloading them to the machine’s main memory takes $180$ms (P50 latency). On this same machine, routing using Equation (2) takes $0.76$ms. As such, so long as a more complex router can achieve the same accuracy as Equation (2) by fetching $16$ fewer shards, but by introducing a computational latency of no greater than roughly $179$ms, the overall efficiency of the query processor improves. As we will show later, our router on the same dataset with the same configuration takes only about $10.1$ms to route queries to shards, leading to a saving of $170.6$ms per query in the given example.
As this example illustrates, a storage-backed, clustering-based ANN search system offers a trade-off space between not just accuracy and speed, but also other efficiency factors such as I/O bandwidth, and memory and cache utilization. This opens the door to more nuanced research. Our work is a step in that direction.
1.4. Existing routers and the role of optimism
Figure 1. (a) Top-$1$ recall vs. percentage of points probed on Text2Image where points have varying norms; (b) Distribution of inner products between a shard and a query on GloVe. Overlaid are scores computed by Mean and NormalizedMean.
As we argued, routing accuracy is increasingly relevant. Surprisingly, with the exception of one recent work that explores supervised learning-to-rank for routing (Vecchiato et al., 2024), the few existing unsupervised routers take the naïve form of Equation (2).
@@ 1.4. Existing routers and the role of optimism
Take Equation (2) as the most prominent example. What we refer to as the Mean router summarizes each shard with its mean point (i.e., $\mu_{i}$ for shard $i$). This is not an unreasonable choice as the mean is the minimizer of the sum of squared deviations from the mean, and is, in fact, natural if KMeans is the clustering algorithm $\zeta$.
@@
(3) $\tau(q)=\operatorname*{arg\,max\quad}^{(\ell)}_{i\in[C]}\langle q,\frac{\mu_{i}}{\lVert\mu_{i}\rVert_{2}}\rangle.$
@@
This formulation is inherited from the familiar Spherical KMeans (Dhillon and Modha, 2001), which is identical to standard KMeans except that, at the end of every iteration, cluster centroids are normalized. Because we can assume that $\lVert q\rVert_{2}=1$ without loss of generality, it is easy to see that Equation (3) routes by the angle between $q$ and the mean vectors.
Intuitively, NormalizedMean seems appropriate for ANN search over a sphere: When norms are constant, all that matters in Equation (1) is the angle between $q$ and data points, rendering it reasonable to rank shards by the angle between their mean and $q$.
@@
Intriguingly, in many circumstances that deviate from that situation, NormalizedMean performs more accurately than the Mean router, as evidenced in Figure LABEL:sub@figure:motivation:mean_vs_normalized_mean. Let us unpack this phenomenon.
@@
Consider a query point $q$ and a single shard $\mathcal{P}$ with mean $\mu$. We visualize in Figure LABEL:sub@figure:motivation:score_distribution the distribution of inner products between $q$ and every point in $\mathcal{P}$. Overlaid is the inner product between $q$ and the mean of points in $\mathcal{P}$, as well as their normalized mean. We observe that $\langle q,\mu/\lVert\mu\rVert_{2}\rangle$ lands in the right tail of the distribution.
@@
That is not surprising for the case where $\lVert u\rVert_{2}=1$ for all $u\in\mathcal{P}$. Clearly $\lVert\mu\rVert_{2}\leq 1$, so that $\langle q,\mu/\lVert\mu\rVert_{2}\rangle=\langle q,\mu\rangle/\lVert\mu\rVert_{2}\geq\langle q,\mu\rangle$; the NormalizedMean router amplifies the Mean router by a factor of $1/\lVert\mu\rVert_{2}$. What is interesting, however, is that the magnitude of this “boost” correlates with the variance of the data within $\mathcal{P}$: The more concentrated $\mathcal{P}$ is around the direction of its mean, the closer the mean is to the surface of the sphere, making $\lVert\mu\rVert_{2}$ larger, so that NormalizedMean applies a smaller amplification to $\langle q,\mu\rangle$. The opposite is true when points are spread out: shards with a higher variance receive a larger lift by the NormalizedMean router!
@@
There are a few caveats. First, it is hard to explain the behavior on point sets with varying norms (i.e., where $\lVert u\rVert_{2}\neq 1$). Second, as we observe in Figure LABEL:sub@figure:motivation:score_distribution, NormalizedMean aggressively overestimates. Nonetheless, the insight that a router’s score for a shard can be influenced by the shard’s variance is worth exploring.
Let us summarize our observation. The estimate given by NormalizedMean paints an optimistic picture of what the maximum inner product between $q$ and points in $\mathcal{P}$ could be. Contrast that with the Mean router, which is naturally a conservative estimate.
@@
We investigate the ramifications of that insight. In particular, the research question we wish to study is whether a more principled approach to designing optimistic routers can lead to more accurate routing decisions on vector sets with variable norms. The nature of this question is not unlike those asked in the online learning literature (Lattimore and Szepesvári, 2020), so it is not surprising that our answer to this question draws from the “principle of optimism in the face of uncertainty.”
1.5. Contributions and outline
@@
We apply the Optimism Principle to routing in clustering-based MIPS. Our formulation, presented in Section 2, rests on the concentration of the inner product distribution between a query $q$ and a set of points in the same shard. In particular, we estimate a score, $\theta_{i}$, for the $i$-th shard such that, with some confidence, the maximum inner product of $q$ with points in that shard is at most $\theta_{i}$. Shards are then ranked by their score for routing. Notice that, when $\theta_{i}=\mu_{i}$, we recover the Mean router, and when $\theta_{i}>\mu_{i}$ routing is optimistic with the confidence parameter determining the degree of optimism.
@@
Building on our proposed formalism of an optimistic router, we outline a general framework that can incorporate as much information as is available about the data distribution to estimate the aforementioned scores. We then present a concrete, assumption-free instance of our algorithm that uses the first and second moments of the empirical inner product distribution only. Furthermore, we make the resulting algorithm space-efficient by designing a sketch (Woodruff, 2014) of the second moment. The end-result is a practical algorithm that is straightforward to implement.
@@
We put our proposal to the test in Section 3 on a variety of ANN benchmark datasets. As our experiments show, our optimistic router achieves the same ANN recall as state-of-the-art routers but with up to a $50\%$ reduction in the total number of data points evaluated per query. We conclude this work in Section 4.
2. Routing by the Optimism Principle: our proposal

### v4

@@ Abstract
Clustering-based nearest neighbor search algorithms partition points into shards to form an index, and search only a subset of shards to process a query. Even though search efficacy is heavily influenced by the algorithm that identifies the shards to probe, it has received little attention in the literature. We study routing in clustering-based maximum inner product search, which includes cosine similarity search. We unpack existing routers and notice the surprising role of optimism. We then take a page from the sequential decision making literature and formalize that insight following the principle of “optimism in the face of uncertainty.” In particular, we present a framework that incorporates the moments of the distribution of inner products within each shard to estimate the maximum inner product. We then develop a practical instance of our algorithm that uses only the first two moments to reach the same accuracy as state-of-the-art routers by probing up to $50\%$ fewer points on benchmark datasets without compromising efficiency. Our algorithm is also space-efficient: we design a sketch of the second moment whose size is independent of the number of points and requires $\mathcal{O}(1)$ vectors per shard.
1 Introduction
A fundamental operation in many applications of machine learning is nearest neighbor search (Bruch, 2024): Given $m$ points denoted $\mathcal{X}\subset\operatorname*{\mathbb{R}}^{d}$, it finds the $k$ closest points to a query $q\in\operatorname*{\mathbb{R}}^{d}$, where closeness is by vector similarity or distance. We focus on inner product, giving the problem of Maximum Inner Product Search (MIPS), which includes Cosine Similarity search as a special case:
$\mathcal{S}=\operatorname*{arg\,max\quad}_{u\in\mathcal{X}}^{(k)}\langle q,u\rangle.$ (1)
@@ 1 Introduction
As $m$ or $d$ grows, an exact solution is often difficult to obtain within a reasonable budget. The problem is thus relaxed to its approximate variant, where we tolerate error to gain speed. The effectiveness of Approximate Nearest Neighbor (ANN) search is characterized by recall or accuracy, defined as the fraction of true nearest neighbors recalled: $\lvert\tilde{\mathcal{S}}\cap\mathcal{S}\rvert/k$, $\tilde{\mathcal{S}}$ being the set returned by the ANN algorithm.
@@
Clustering-based ANN search: ANN algorithms come in various flavors, from trees (Bentley, 1975; Dasgupta and Sinha, 2015), LSH (Indyk and Motwani, 1998), to graphs (Malkov and Yashunin, 2020; Jayaram Subramanya et al., 2019). Refer to (Bruch, 2024) for a review of this subject. The method relevant to us is the clustering-based approach, also known as Inverted File, which has received much attention in the literature (Auvolat et al., 2015; Babenko and Lempitsky, 2012; Chierichetti et al., 2007; Bruch et al., 2024; Douze et al., 2024) and is widely adopted in industrial applications. [^1] See, for example, https://turbopuffer.com/blog/turbopuffer, https://www.pinecone.io/blog/serverless-architecture/#Pinecone-serverless-architecture, and https://research.google/blog/announcing-scann-efficient-vector-similarity-search/.
In this paradigm, $\mathcal{X}$ is partitioned into $C$ shards using a clustering function $\zeta:\operatorname*{\mathbb{R}}^{d}\rightarrow[C]$—typically KMeans with $C=\mathcal{O}(\sqrt{m})$. Accompanying this index is a router $\tau:\mathbb{R}^{d}\rightarrow[C]^{\ell}$, which returns $\ell$ shards likely to contain $q$’s nearest neighbors. If $\mu_{i}$ is the mean of the $i$-th shard, a common router is:
$\tau(q)=\operatorname*{arg\,max\quad}^{(\ell)}_{i\in[C]}\langle q,\mu_{i}\rangle.$ (2)
@@
Processing a query $q$ involves two independent subroutines. The first, “routing,” returns $\ell$ shards using $\tau(q)$, and the subsequent step, “scoring,” computes inner products between $q$ and the points in the $\ell$ shards. While the literature has focused on scoring (Jégou et al., 2011; Ge et al., 2014; Wu et al., 2017; Andre et al., 2021; Kalantidis and Avrithis, 2014; Johnson et al., 2021; Norouzi and Fleet, 2013), routing has received little attention. This work turns squarely to the routing step.
@@
We emphasize that, routing is independent of the scoring algorithm. For example, we may route using Equation (2) or our novel router to be introduced later. This choice has no algorithmic bearing on scoring whatsoever. In other words, scoring can be accomplished by a linear scan of the chosen shards, computing inner products with Product Quantization (PQ) (Jégou et al., 2011), or searching each shard using a graph ANN algorithm Jayaram Subramanya et al. (2019); Malkov and Yashunin (2020). All these choices are independent of the type of router utilized in the earlier stage.
@@
The importance of routing: The historical focus on scoring makes sense. Even though routing narrows down the search space, selected shards may contain a large number of points. Scoring must thus be efficient and effective.
@@
We argue that the routing step is important in its own right. First, a more accurate [^2] We quantify a router’s accuracy by the ANN recall in a setup where scoring is by a linear scan over the identified shards, so that the only source of error is the routing procedure. router leads to fewer data points being scored. For example, if shards are balanced in size, access to an oracle router (i.e., one that identifies the shard with the true nearest neighbor) means that the scoring stage must only compute inner products for $m/C$ points to find the top-$1$ point.
@@
The second reason pertains to scale. As size and dimensionality grows, it is often infeasible to keep the entire index in memory, compression techniques such as PQ notwithstanding. Much of the index must instead rest on secondary storage—cheap but high-latency storage such as disk or blob storage—and accessed when necessary. That line of reasoning has led to the emergence of disk-based graph indexes (Jayaram Subramanya et al., 2019; Singh et al., 2021; Jaiswal et al., 2022).
@@
Translating the same rationale to the clustering-based paradigm implies that shards rest outside of memory, and when a router identifies a subset of shards, scoring must first fetch those shards from storage. This paradigm has gained traction in real-world database systems.333c.f., documents linked in the footnote above. This is the context in which we present our research: Storage-backed, clustering-based ANN search systems.
@@
In this framework, a more accurate router lowers the volume of data transferred between storage and memory with implications for storage bandwidth and in-memory cache space utilization, and retrieval throughput. We have included an illustrative example in Appendix A to support this statement.
@@
Existing routers and the role of optimism: We review the relevant literature on routers in Appendix B. Among existing routers, Equation (2), which we call Mean, is the most prominent. It summarizes each shard by its mean point ($\mu_{i}$’s). This is not an unreasonable choice as the mean is the minimizer of the sum of squared deviations from the mean, and is, in fact, natural if $\zeta$ is KMeans.
Figure 1: (a) Top-11 recall vs. percentage of points probed on Text2Image where points have varying norms; (b) and (c) Distribution of inner products between a shard and a query on GloVe. Overlaid are scores computed by Mean and NormalizedMean.
@@
$\tau(q)=\operatorname*{arg\,max\quad}^{(\ell)}_{i\in[C]}\langle q,\frac{\mu_{i}}{\lVert\mu_{i}\rVert_{2}}\rangle.$ (3)
@@
This formulation is inherited from spherical KMeans (Dhillon and Modha, 2001), which is identical to standard KMeans except that cluster centroids are normalized after every iteration. Because we can assume that $\lVert q\rVert_{2}=1$ without loss of generality, Equation (3) routes by the angle between $q$ and $\mu_{i}$’s.
@@
NormalizedMean seems appropriate for search over a sphere: When norms are constant, only angles matter in Equation (1), making it reasonable to rank shards by the angle between means and $q$. Intriguingly, in many circumstances that deviate from that situation, NormalizedMean performs more accurately than Mean, as evidenced in Figure LABEL:sub@figure:motivation:mean_vs_normalized_mean. Let us unpack this phenomenon.
@@
Consider $q$ and a single shard $\mathcal{P}$ with mean $\mu$. Figure LABEL:sub@figure:motivation:score_distribution visualizes the distribution of inner products of $q$ and all points in $\mathcal{P}$. Overlaid is the inner product of $q$ and $\mu$, as well as their normalized mean. We observe that $\langle q,\mu/\lVert\mu\rVert_{2}\rangle$ lands in the right tail of the distribution.
@@
That is not surprising when $\lVert u\rVert_{2}=1$ for all $u\in\mathcal{P}$. Clearly $\lVert\mu\rVert_{2}\leq 1$, so that $\langle q,\mu/\lVert\mu\rVert_{2}\rangle=\langle q,\mu\rangle/\lVert\mu\rVert_{2}\geq\langle q,\mu\rangle$; NormalizedMean amplifies Mean by a factor of $1/\lVert\mu\rVert_{2}$. What is interesting, however, is that the magnitude of this “boost” correlates with the variance of $\mathcal{P}$: The more concentrated $\mathcal{P}$ is around the direction of $\mu$, the closer $\mu$ is to the surface of the sphere, making $\lVert\mu\rVert_{2}$ larger, so that NormalizedMean applies a smaller boost to $\langle q,\mu\rangle$. The opposite is true when points are spread out: shards with a higher variance receive a larger lift by NormalizedMean.
@@
There are a few caveats. First, it is hard to explain the behavior on sets with varying norms. Second, as we observe in Figure LABEL:sub@figure:motivation:score_distribution_overzealous, NormalizedMean’s estimate can be overly aggressive. Nonetheless, the insight that a router’s score for a shard can be influenced by the shard’s variance is worth exploring.
Contributions and outline: From the exercise above, we learned that NormalizedMean paints an optimistic picture of what the maximum inner product of $q$ and $\mathcal{P}$ could be. Mean, however, gives a conservative estimate. We investigate the ramifications of that insight in this work.
@@
In particular, the research question we study is whether a more principled approach to designing optimistic routers leads to more accurate routing on vector sets with variable norms. This question is not unlike those asked in the online learning literature (Lattimore and Szepesvári, 2020), so it is not surprising that our answer draws from the “principle of optimism in the face of uncertainty.”
@@
We apply the Optimism Principle to routing in clustering-based MIPS. Our method, presented in Section 2, uses the concentration of inner products between a query $q$ and points in a shard. In particular, we estimate a score, $\theta_{i}$, for the $i$-th shard such that, with some confidence, the maximum inner product of $q$ with points in that shard is at most $\theta_{i}$. Shards are then ranked by this score. When $\theta_{i}=\mu_{i}$, we recover the Mean router, and when $\theta_{i}>\mu_{i}$ routing is optimistic with a controllable degree of optimism. Importantly, our routing function is independent of how shards are formed.
@@
We outline a general framework that can incorporate as much information as is available about the data distribution to estimate $\theta_{i}$’s. We then present a concrete, assumption-free instance of our algorithm that uses the first and second moments of the empirical inner product distribution only. Furthermore, we make the resulting algorithm space-efficient by designing a sketch (Woodruff, 2014) of the second moment. The end-result is a practical algorithm that is straightforward to implement.
@@
We test our proposal in Section 3 on a variety of ANN benchmark datasets. As our experiments show, our optimistic router achieves the same recall as state-of-the-art routers but with up to a $50\%$ reduction in the total volume of points evaluated per query. We conclude this work in Section 4.
2 Routing by the Optimism Principle: our proposal